OpenStudy (anonymous):

Question 5.5. You have 0.5 L of air at 203 K in an expandable container that allows for a constant pressure. You heat the container to 373 K. What is the volume of air? (Points : 1) 0.27 L 0.67 L 0.92 L 2.7 L

OpenStudy (anonymous):

@carsonsi

OpenStudy (anonymous):

0.67 L i believe

OpenStudy (aaronq):

\(\sf PV=nRT\) P, R, n constant \(\sf \cancel PV=\cancel{nR}T\rightarrow \dfrac{V}{T}=constant\) therefore \(\sf \dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\)